# Size-structured populations with growth fluctuations: Feynman–Kac formula and decoupling

### v1

@@ 3.2 Lineage representation of population statistics
In addition to the decoupling results, we describe the relationship between lineage and population observables via path integral representations. These representations are also known as the “Many-to-one” formula, as they relate the population process (the many body theory in physics terminology) to the lineage process [35, 34]. For strong decoupling, we can obtain expectations with respect to the population distribution by averaging, yielding
$\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]\sim\frac{\mathbb{E}_{\rm l}^{\rm path}[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}]}{\mathbb{E}_{\rm l}^{\rm path}[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}]}.$ (19)
where $\mathbb{E}_{\rm l}^{\rm path}$ is used to represent averages with respect to the lineage sample path distribution on $[0,t)$ (rather than over $\nu({\boldsymbol{x}},t)$).
Additionally, we will show that when division is symmetric, but the internal variable and size phenotype do not necessarily decouple, the right-hand side of equation (19) has an interpretation as an average with respect to a mass-weighted distribution (section 6).
@@ 4.1 Size-only process
A special case of the model described above is the size-only process obtained by setting $\lambda({\boldsymbol{x}})=\lambda_{0}$. In this case, the time-homogeneous solution to equation (6), $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ factors out into a growth distribution $\nu({\boldsymbol{x}})$ and a size distribution $w({\boldsymbol{y}})$. The latter must now satisfy
$\partial_{y_{c}}w_{\rm l}(y_{c},y_{b})=-\varphi(y_{c},y_{b})w_{\rm l}(y_{c},y_{b}).$ (20)
@@
$w_{\rm l}^{b}(y)=Z^{-1}w_{\rm l}(y,y),$ (22)
where the normalization constant $Z=\int_{-\infty}^{\infty}w_{\rm l}(y,y)dy$ ensures this is a density in $y$. Using the boundary conditions given in equation (8), we find that $\omega^{b}_{\rm l}$ satisfies the integral equation
$w_{\rm l}^{b}(y)=\int_{-\infty}^{\infty}\int_{y_{b}^{\prime}}^{\infty}u(y|y_{c}^{\prime})f(y_{c}^{\prime}|y_{b}^{\prime})w_{\rm l}^{b}(y_{b}^{\prime})\,dy_{c}^{\prime}dy_{b}^{\prime},$ (23)
@@
In the special case of symmetric division, where $u(y_{b}|y_{c})=\delta(y_{b}-y_{c}+\ln(2))$, one can check that for any solution $w_{\rm l}(y_{c},y_{b})$ to the lineage size distribution equations (20) and (23), the lineage distribution divided by cell mass $w_{\rm l}(y_{c},y_{b})e^{-y_{c}}$ must satisfy the population distribution equations (26) and (27). We therefore have the proportionality $w_{\rm p}(y_{c},y_{b})\propto w_{\rm l}(y_{c},y_{b})e^{-y_{c}}$. In A, we discuss the cell size distributions marginalized over birth size and show how they reduce to known forms that appear in other literature.
@@
where $\phi$ and $\Phi$ are, respectively, the pdf and cdf of standard Gaussian distributions. Then the distribution of the final size conditioned on the initial size is approximately ${\rm Normal}(y_{b}(1-\alpha)+\ln(2),\sigma_{Y})$, provided $\sigma_{Y}\ll 1$. This is the commonly used linear regression model for size dynamics [20, 10, 1]. The approximation comes from the fact that $f(y_{c}|y_{b})$ is supported on $y_{c}\geq y_{b}$, whereas in the exactly Gaussian model $y_{c}$ may be less than $y_{b}$. We include a brief discussion of this autoregressive model and the biological significance of $\alpha$ in B. Note that if we take $u(\cdot|y_{c}^{\prime})$ to be Gaussian with mean $y_{c}^{\prime}$ and variance $\sigma_{d}^{2}$, then we can absorb the variance introduced by division into the parameter $\sigma_{Y}^{2}$, thus we can set $u(y|y_{c}^{\prime})=\delta_{y_{c}^{\prime}-\ln(2)}$, without loss of generality. As long as $\alpha\in(0,2)$, the size distribution will remain stable.
Figure 2: (left) Simulations of three models illustrating examples of (top to bottom) (SD), (WDl), and no decoupling. For all models $d=1$ and the growth rate function is $\lambda=1+x$. For (SD) we have simulated an OU process for the $\boldsymbol{X}$} dynamics with $\theta=1$, $\sigma^{2}=0.01$, $\alpha=1/2$, $\sigma_{Y}=0.01$. For (WDl) we have simulated a model with $\mathcal{L}=0$ and $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})$ is Gaussian with mean 0 and standard deviation $0.1$. The last model combines the OU dynamics with this division kernel. (middle) The convergence of the empirical correlation coefficients between size variables and $X^{2}$ in the lineage and (right) population distribution.
4.2 OU Growth rate dynamics (strong decoupling)
As an example where the growth rate fluctuates, consider an Ornstein-Uhlenbeck (OU) process for the growth dynamics within the cell-cycle, which has previously been studied in [48, 20, 29, 49, 11, 9, 27]. For simplicity we take $d=1$ (the multivariate version of this process is discussed in C). The process $X(t)$ is given by
@@ 4.2 OU Growth rate dynamics (strong decoupling)
The stationary density (which is the normalized Perron eigenvector of $\mathcal{L}$) of the size-independent process is a ${\rm Normal}(0,\sigma^{2}/\theta)$.
We consider this model division rate $\beta=\lambda\varphi$, where $\lambda(x)=\lambda_{0}+x$ and $\varphi$ is given by equation (28). Although $\lambda$ is not strictly positive, we work in the small noise regime where the negative growth rates are exponentially unlikely, and therefore this can be considered as an approximation to $(\lambda_{0}+x)1_{x>-\lambda_{0}}$. As was shown in [20], if $r(x|x^{\prime})=\delta(x-x^{\prime})$ (implying SD), $X(t)$ and $Y(t)$ are independent in both the lineage and population ensemble. Note that technically, because $\lambda$ is not positive, we can obtain non-physical solutions where the growth rate becomes negative, but these have a negligible effect in the small noise regime. Figure 2 (top row) illustrates this numerically. Here, we have computed the empirical correlation coefficients $\widehat{{\rm corr}}_{T}(X^{2},Y_{c})$, $\widehat{{\rm corr}}_{T}(X^{2},Y_{c}-y_{b})$ computed from averages over a time window $T$. These show a clear convergence to zero as we looked over longer time windows (lineage dynamics) and larger population sizes (population dynamics).
@@
We computed these correlations specifically (rather than correlations with $X$) because $X$ did not appear to have detectable correlations in the population distribution of the second model, and therefore did not illustrate the lack of decoupling in the population ensemble.
4.3 Division noise
We add noise in $X$ at division to the OU model by taking $r(\cdot|X^{\prime})$ to be a ${\rm Normal}(\mu_{X}+c\,x^{\prime},\sigma^{2}_{X})$ density. Taking the $\sigma^{2}/\theta\to 0$ gives a model with constant growth over the cell-cycle ($\mathcal{L}=0$). This is known as the random growth rate model or the RG model for short. Simulations of this model are also shown in figure 2. The RG falls into the case of weak decoupling (WDl) if we use the same model for $\beta$ as the previous subsection. Consistent with our results described above, the internal variable and size phenotype decouple in the lineage, but not the population ensemble. When both sources of noise are present, without some fine-tuning of the division distribution, there is no decoupling, as can be seen in the bottom row of figure 2.
@@ 5 Strong decoupling and Feynman–Kac formula
In this section, we discuss the eigenvalue problem in equation (16) from the perspective of the Feynman–Kac formula. We do not go into the technical details of the Feynman–Kac theory. The general theory of Feynman–Kac semigroups can be found in [36], while applications to structured population models are discussed in [35, 6].
5.1 Time-changed process
@@
The marginal size process generated by the model described in section 4 is then given by
@@
Assuming $\lambda$ is positive, $T(t)$ is invertible and we can write $\widetilde{{\boldsymbol{Y}}}(s)={\boldsymbol{Y}}(T^{-1}(s))$ This is because each time increment in the stochastic time variable corresponds to an increment $\lambda({\boldsymbol{X}}_{t})dt$ in the original time variable.
To derive equation (32), consider the probability $\widetilde{{\boldsymbol{Y}}}(t)$ dividing (meaning it is resampled from $u$) in an interval $[T,T+dT)$, which by definition of $\varphi$ is
@@
The equivalent probability for ${\boldsymbol{Y}}(t)$ is
$\begin{split}&\mathbb{P}\left(\text{${\boldsymbol{Y}}(t)$ divides in $[t,t+dt)$}\Big{|}{\boldsymbol{Y}}(t),{\boldsymbol{X}}(t)\right)\\ &=\mathbb{P}\left(\text{$\widetilde{{\boldsymbol{Y}}}(T)$ divides in $[T(t),T(t)+T^{\prime}(t)dt)$}\Big{|}\widetilde{{\boldsymbol{Y}}}(T),{\boldsymbol{X}}(t)\right)\\ &=\lambda({\boldsymbol{X}}(t))\varphi({\boldsymbol{Y}}(t))dt.\end{split}$ (34)
This transformation procedure is shown in figure 3 (A). In figure 3 (B), we have shown trajectories of a process with OU growth rate. Note that the time change can be performed even when ${\boldsymbol{X}}(t)$ depends on ${\boldsymbol{Y}}(t)$, because by the definition of $T(t)$, $T^{\prime}(t)|{\boldsymbol{X}}(t)$ is always deterministically equal to $\lambda({\boldsymbol{X}}(t))$.
@@
5.2 The Feynman–Kac formula
When (SD) holds, the population distribution may be obtained via expectations involving $e^{T(t)}$. Recall that in this case $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})=\delta({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})$ and $\mathcal{L}$ generates a stationary ergodic Markov process. Let
$g_{\rm p}^{f}({\boldsymbol{x}},t)=\mathbb{E}_{\rm l}^{\rm path}\left[f({\boldsymbol{X}}(t))e^{T(t)}\Big{|}{\boldsymbol{X}}(0)={\boldsymbol{x}}\right].$ (35)
@@ 5.2 The Feynman–Kac formula
and $g_{\rm p}^{f}({\boldsymbol{x}},0)=f({\boldsymbol{x}})$. Equation (36) can be derived for SDEs using Itô’s lemma (see [38]), but it can also be derived for a more general process by looking directly at the forward variation of $g_{\rm p}^{f}$ and Taylor expanding, similar to textbook derivations of the Kolmogorov backward equations [17].
A consequence of the Krein–Rutman theorem (a generalization of the Perron-Frobenius theorem for operators [43]) is that the non–conservative operator $\hat{\mathcal{L}}_{\lambda}$ has a dominant eigenvalue $\Lambda$ corresponding to a strictly positive, normalized right eigenfunction $\nu_{\rm p}$ (the population density); that is,
@@
$\frac{g_{\rm p}^{f}({\boldsymbol{x}},t)}{g_{\rm p}^{1}({\boldsymbol{x}},t)}=\frac{\mathbb{E}_{\rm l}^{\rm path}[f({\boldsymbol{X}}(t))e^{T(t)}]}{\mathbb{E}_{\rm l}^{\rm path}[e^{T(t)}]}$ (42)
is the ratio appearing in equation (19). Therefore, as claimed, the population-level expectations can be obtained as the $t\to\infty$ limit of weighted lineage expectations.
@@
$\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]-\mathbb{E}_{\rm l}[f({\boldsymbol{X}}(t))]={\rm Cov}_{\rm l}^{\rm path}\left(f({\boldsymbol{X}}(t)),\frac{e^{T(t)}}{\mathbb{E}_{\rm l}\left[e^{T(t)}\right]}\right).$ (43)
@@ 5.3 Numerical example
The estimator $\hat{f}_{p}(t)$ of $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]$ is
@@
Intuitively, we expect this to give a good approximation to $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]$ when both $t$ and $m$ are sufficiently large; however, as shown in refs [26, 19, 42], estimating expectations of this form is difficult in practice. In particular, it is shown that there are two sources of error, a finite time bias and a sampling bias, the interplay of which leads to non-monotonic convergence in $t$. These works focus on the estimation of the dominant growth rate $\Lambda=\lim_{t\to\infty}t^{-1}\ln\mathbb{E}_{\rm l}[e^{T(t)}]$, rather than the expectations with respect to $\nu_{\rm p}$, but we expect $\hat{f}_{\rm p}(t)$ to be plagued by the same issues. Nonetheless, we have found that the Feynman–Kac estimator does provide some information about the selection bias as shown in figure 4. For any given time value, the estimator is too noisy to detect differences between the lineage and population expectation, yet the time-average gives good agreement. We leave it to future work to explore the practical issues surrounding these estimates.
Figure 4: The estimator ($\hat{f}_{p}$) as a function of the time-window $t$ (solid blue line), its time average (dashed blue line), and the theoretical value of the population expectation. Remember that in this model, the lineage expectation is 0.

### v2

@@ 3.2 Lineage representation of population statistics
In addition to the decoupling results, we describe the relationship between lineage and population observables via path integral representations. These representations are also known as the “Many-to-one” formula, as they relate the population process (the many-body theory in physics terminology) to the lineage process [48, 47]. For strong decoupling, we can obtain expectations with respect to the population distribution by averaging, yielding
$\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]\sim\frac{\mathbb{E}_{\ell}^{\rm path}[f({\boldsymbol{X}}(t))e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}]}{\mathbb{E}_{\ell}^{\rm path}[e^{\int_{0}^{t}\lambda({\boldsymbol{X}}(s))ds}]}.$ (20)
where $\mathbb{E}_{\ell}^{\rm path}$ is used to represent averages with respect to the lineage sample path distribution on $[0,t)$ (rather than over $\nu({\boldsymbol{x}},t)$). Additionally, we will show that when division is symmetric, but the internal variable and size phenotype do not necessarily decouple, the right-hand side of equation (20) has an interpretation as an average with respect to a mass-weighted distribution (section 6).
@@ 4.1 Size-only process
A special case of the model described above is the size-only process obtained by setting ${\boldsymbol{X}}={\boldsymbol{x}}_{0}$ (by letting $\mathcal{L}=0$ and assuming SD) and hence $\lambda({\boldsymbol{x}}_{0})=\lambda_{0}$. In this case, the phenotype of a cell is defined only by the size variable. The time-homogeneous solution to equation (7), $\rho({\boldsymbol{x}},{\boldsymbol{y}})$ factors out into a growth distribution $\nu({\boldsymbol{x}})$ and a size distribution $w({\boldsymbol{y}})$. The lineage density of the size variable $w_{\ell}({\boldsymbol{y}})=w_{\ell}(y_{c},y_{b})$ is given by
$\partial_{y_{c}}w_{\ell}(y_{c},y_{b})=-\varphi(y_{c},y_{b})w_{\ell}(y_{c},y_{b}).$ (21)
@@
$w_{\ell}^{b}(y)=Z^{-1}w_{\ell}(y,y),$ (23)
where the normalization constant $Z=\int_{-\infty}^{\infty}w_{\ell}(y,y)dy$ ensures this is a density in $y$. Using the boundary conditions given in equation (10), we find that $w^{b}_{\ell}$ satisfies the integral equation
$w_{\ell}^{b}(y)=\int_{-\infty}^{\infty}\int_{y_{b}^{\prime}}^{\infty}u(y|y_{c}^{\prime})f(y_{c}^{\prime}|y_{b}^{\prime})w_{\ell}^{b}(y_{b}^{\prime})\,dy_{c}^{\prime}dy_{b}^{\prime},$ (24)
@@
In the special case of symmetric division, where $u(y_{b}|y_{c})=\delta(y_{b}-y_{c}+\ln(2))$, one can check that for any solution $w_{\ell}(y_{c},y_{b})$ to the lineage size distribution equations (21) and (24), the lineage distribution divided by cell mass $w_{\ell}(y_{c},y_{b})e^{-y_{c}}$ must satisfy the population distribution equations (27) and (28). We therefore have the proportionality $w_{\rm p}(y_{c},y_{b})\propto w_{\ell}(y_{c},y_{b})e^{-y_{c}}$. In A, we discuss the cell size distributions marginalized over birth size and show how they reduce to known forms that appear in other literature.
@@
where $\phi$ and $\Phi$ are, respectively, the pdf and cdf of standard Gaussian distributions. The distribution of the final size conditioned on the initial size is approximately ${\rm Normal}(y_{b}(1-\alpha)+\ln(2),\sigma_{Y})$, provided $\sigma_{Y}\ll 1$. This is the commonly used linear regression model for size dynamics [26, 11, 1]. The approximation comes from the fact that $f(y_{c}|y_{b})$ is supported on $y_{c}\geq y_{b}$, whereas in the exactly Gaussian model $y_{c}$ may be less than $y_{b}$. We include a brief discussion of this autoregressive model and the biological significance of $\alpha$ in B. Note that if we take $u(\cdot|y_{c}^{\prime})$ to be Gaussian with mean $y_{c}^{\prime}$ and variance $\sigma_{d}^{2}$, then we can absorb the variance introduced by division into the parameter $\sigma_{Y}^{2}$, thus we can set $u(y|y_{c}^{\prime})=\delta_{y_{c}^{\prime}-\ln(2)}$, without loss of generality. As long as $\alpha\in(0,2)$, the size distribution will remain stable.
Figure 2: (left) Simulations of three models illustrating examples of (top to bottom) (SD), (WD), and no decoupling. For all models $d=1$ and the growth rate function is $\lambda=1+x$. For (SD) we have simulated an OU process for the $\boldsymbol{X}$} dynamics with $\theta=1$, $\sigma^{2}=0.01$, $\alpha=1/2$, $\sigma_{Y}=0.01$. For (WD) we have simulated a model with $\mathcal{L}=0$ and $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})$ is Normal distribution with mean 0 and standard deviation $\sigma_{x}=0.1$. The last model combines the OU dynamics with this division kernel. (middle) The convergence of the empirical correlation coefficients between size variables and $X$ in the lineage and (right) population distribution. The dashed lines show the running average of the correlations and the shaded area shows $\pm$ one standard deviation.
4.2 OU and RG Growth rate dynamics
As an example where the growth rate fluctuates, consider an Ornstein-Uhlenbeck (OU) process for the growth dynamics within the cell-cycle, which has previously been studied in references [62, 26, 42, 65, 13, 10, 40]. For simplicity we take $d=1$ (the multivariate version of this process is discussed in C). The process $X(t)$ is given by
@@ 4.2 OU and RG Growth rate dynamics
The stationary density (which is the normalized Perron eigenvector of $\mathcal{L}$) of the size-independent process is a ${\rm Normal}(0,\sigma^{2}/(2\theta))$.
We consider this model division rate $\beta=\lambda\varphi$, where $\lambda(x)=\lambda_{0}+x$ and $\varphi$ is given by equation (29). Although $\lambda$ is not strictly positive, we work in the small noise regime where the negative growth rates are exponentially unlikely; therefore, this can be considered an approximation to $(\lambda_{0}+x)1_{x>-\lambda_{0}}$. As was shown in [26], if $r(x|x^{\prime})=\delta(x-x^{\prime})$ (which together with equation (31) implies SD), $X(t)$ and $Y(t)$ are independent in both the lineage and population ensembles. Note that technically, because $\lambda$ is not positive, we can obtain non-physical solutions where the growth rate becomes negative; however, but these have a negligible effect in the small noise regime. Figure 2 (top row) illustrates this numerically. Here, we have computed the empirical correlation coefficients $\widehat{{\rm corr}}_{T}(X,Y_{c})$. These show a clear convergence to zero as we looked over longer time windows (lineage dynamics) and larger population sizes (population dynamics).
@@
We contrast the OU model with the so-called Random Growth rate (RG) Model where $\mathcal{L}=0$ and $r(\cdot|X^{\prime})$ is taken to be a ${\rm Normal}(0,\sigma_{X})$ density. Simulations of this model are also shown in figure 2. The RG falls into the case of weak decoupling (WD) if we use the same model for $\beta$ as in the previous subsection. Consistent with our results described above, the internal variable and size phenotype decouple in the lineage, but not in the population ensemble.Finally, we simulated the OU model but violated condition $1$ by setting $\beta=\lambda_{0}\varphi$, so there is not $\boldsymbol{x}$} dependence on the division rate. As can be seen in the bottom row of figure 2, the correlations between $X$ and $Y_{c}$ persist as $t\to\infty$ both in the lineage and population.
@@ 5 Strong decoupling and Feynman–Kac formula
In this section, we discuss the eigenvalue problem in equation (17) from the perspective of the Feynman–Kac formula. We do not go into the technical details of the Feynman–Kac theory, which can be found in [49]. Applications to structured population models are discussed in [48, 7].
5.1 Time-changed process and Feynman–Kac formula
@@
The marginal size process generated by the model described in section 4.1 is then given by
@@
Assuming $\lambda$ is positive, $T(t)$ is invertible, and we can write $\widetilde{{\boldsymbol{Y}}}(s)={\boldsymbol{Y}}(T^{-1}(s))$ This is because each time increment in the stochastic time variable corresponds to an increment $\lambda({\boldsymbol{X}}_{t})dt$ in the original time variable.
To derive equation (33), consider the probability $\widetilde{{\boldsymbol{Y}}}(t)$ dividing (meaning it is resampled from $u$) in an interval $[T,T+dT)$, which, by the definition of $\varphi$, is
@@
Hence, the probability for ${\boldsymbol{Y}}(t)$ to divide in $[t,t+dt)$ is $\lambda({\boldsymbol{X}}(t))\varphi({\boldsymbol{Y}}(t))dt$. This transformation procedure is shown in figure 3 (A). In figure 3 (B), we have shown trajectories of a process with an OU growth rate. Note that the time change can be performed even when ${\boldsymbol{X}}(t)$ depends on ${\boldsymbol{Y}}(t)$, because by the definition of $T(t)$, $T^{\prime}(t)|{\boldsymbol{X}}(t)$ is always deterministically equal to $\lambda({\boldsymbol{X}}(t))$.
@@
When (SD) holds, the population distribution may be obtained via expectations involving $e^{T(t)}$. Recall that in this case $r({\boldsymbol{x}}|{\boldsymbol{x}}^{\prime})=\delta({\boldsymbol{x}}-{\boldsymbol{x}}^{\prime})$ and $\mathcal{L}$ generate a stationary ergodic Markov process. Let
$g_{\rm p}^{f}({\boldsymbol{x}},t)=\mathbb{E}_{\ell}^{\rm path}\left[f({\boldsymbol{X}}(t))e^{T(t)}\Big|{\boldsymbol{X}}(0)={\boldsymbol{x}}\right].$ (35)
@@
and $g_{\rm p}^{f}({\boldsymbol{x}},0)=f({\boldsymbol{x}})$. Equation (36) can be derived for SDEs using Itô’s lemma (see [52]), but it can also be derived for a more general process by looking directly at the forward variation of $g_{\rm p}^{f}$ and performing a Taylor expansion, similar to the textbook derivations of the Kolmogorov backward equations [20].
It is known that (see [49]) the operator $\hat{\mathcal{L}}_{\lambda}$ has a dominant eigenvalue $\Lambda$ corresponding to a strictly positive, normalized right eigenfunction $\nu_{\rm p}$ (the population density); that is,
@@
$\frac{g_{\rm p}^{f}({\boldsymbol{x}},t)}{g_{\rm p}^{1}({\boldsymbol{x}},t)}=\frac{\mathbb{E}_{\ell}^{\rm path}[f({\boldsymbol{X}}(t))e^{T(t)}]}{\mathbb{E}_{\ell}^{\rm path}[e^{T(t)}]}$ (42)
The ratio appears in equation (20). Therefore, as claimed, the population-level expectations can be obtained as the $t\to\infty$ limit of weighted lineage expectations.
@@
$\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]-\mathbb{E}_{\ell}[f({\boldsymbol{X}}(t))]={\rm Cov}_{\ell}^{\rm path}\left(f({\boldsymbol{X}}(t)),\frac{e^{T(t)}}{\mathbb{E}_{\ell}\left[e^{T(t)}\right]}\right).$ (43)
@@
Figure 4: (left) Estimated phenotypic average $\hat{f}_{p}(t)$ compared to theoretical expectations $E_{\rm p}[X]$ and $\mathbb{E}_{\ell}[X]$, with cumulative time average shown in blue. (right) Error showing deviation of time-averaged estimate from true value and standard error. Results from $m=100$ independent lineage simulations of OU with $T=1000$ with $\sigma^{2}=0.01$, $\theta=1$ and $\sigma_{Y}=0.05$.
@@ 5.2 Numerical example
For the $m$ lineages, we compute the estimate
@@
of $\mathbb{E}_{\rm p}[f({\boldsymbol{X}}(t))]$. In figure 4, we have plotted this estimator as a function of time for $m=500$ lineages. We also plot the time-average $\langle\hat{f}_{p}\rangle_{t}=\frac{1}{t}\int_{0}^{t}\hat{f}_{p}(s)ds$ and the running standard deviation. We observe that $\langle\hat{f}_{p}\rangle_{t}$ tends towards the true population average, although the variance of this estimator diverges.
To explain the behavior in Figure 4, note that the Law of Large Numbers ensures that for fixed $t$, as $m\to\infty$, $\hat{f}_{p}(t)\to\mathbb{E}_{\ell}[f(X(t))\,e^{T(t)}]/\mathbb{E}_{\ell}[e^{T(t)}]$. The equivalence between this lineage-based quantity and the population expectation holds only asymptotically in time; thus, so the lineage must be run for a burn-in period. However, the variance of $e^{T_{i}(t)}$ grows exponentially in $t$, which implies that longer burn-in periods yield larger standard errors. Consequently, accurate estimation of the population expectation requires, in principle, exponentially many lineages. The exact scaling between time and the required number of lineages depends on the model and the observable. Nevertheless, in the present example, the estimator performs well even over time scales that can realistically be sampled in a mother-machine experiment.
A similar situation is considered in the context of growth rate estimation in references [39, 24, 57], In particular, it is shown that there are two sources of error, a finite time bias and a sampling bias, the interplay of which leads to non-monotonic convergence in $t$. These works focus on the estimation of the dominant growth rate $\Lambda=\lim_{t\to\infty}t^{-1}\ln\mathbb{E}_{\ell}[e^{T(t)}]$, rather than the expectations with respect to $\nu_{\rm p}$, but we expect $\hat{f}_{\rm p}(t)$ to be plagued by the same issues.